OpenStudy (anonymous):

At what angle (to the horizontal) must the projectile be fired to hit a target positioned at a range of 10.0 km away if the initial speed of the projectile v0 = 500 m/s? I know I must find both the horizontal and vertical velocities, then use arctan, but have no idea how to do the horizontal velocity.

OpenStudy (anonymous):

Horizontal component of the velocity is Vx=500*Cos(alpha) you can find vertical component on your own. 

OpenStudy (anonymous):

I'm still not sure of the steps to get the answer. The answer should be 11.8 degrees. The vertical velocity would be Vy = 500(sin)alpha. What do I do with that after? Someone said the equation should be this: Range = u^2Sin2θ/g 10*10^3 = 500^2*Sin2(θ)/9.8 = 11.54 degree *Does anyone know how he got 10*10^3?

OpenStudy (wolfe8):

10*10^3 is the range you want, that is 10 km written meters since the velocity given is in m/s.

OpenStudy (anonymous):

Oh duh! thank you

OpenStudy (wolfe8):

You're welcome. It's helpful to convert everything into one unit you are comfortable working with before you start solving. It helps reduce mistake :)

OpenStudy (anonymous):

Thanks again, will try to remember that from now on. I try but keep forgetting. :)

OpenStudy (anonymous):

I am having trouble getting the right answer. It should be 11.54 degrees, I think I am using my calculator wrong or calculating wrong. After plugging everything in I get: Theta = arcsin 1/2* (10*10^3 (20)/500^2) Theta = arcsin .4 = 23.6 degrees. I am in degree mode, can't figure out what I am doing wrong?

OpenStudy (wolfe8):

This is what I get when I solve for theta. \[\theta =\frac{ 1 }{ 2 } \sin^{-1} \frac{(10X10^{3})9.8 }{ 500^{2} }\] Make sure you solve everything in brackets first, then multiply, divide, then arsin it. Then divide by 2

OpenStudy (anonymous):

Ah I think I was maybe maybe using the wrong formula. The one my professor gave me was Range = \[Range = V0^2\sin^2\alpha/2g\] However I found on wikipedia the formula should be \[\alpha = 1/2 \sin^{-1} (gd/v0^2)\], g being =10 (my prof allows us to use 10 for simplicity) and d = distance So why 2g vs. g?

OpenStudy (wolfe8):

Sorry I have to go now. Good luck.

OpenStudy (anonymous):

So with g = 10 it is 11.8 degrees. However I'm not sure I understand why his formula was different. Thanks though!